Circular No. 6073

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS R. M. Hjellming and M. Rupen, National Radio Astronomy Obser-

vatory, report the results of Very Large Array radio observations of GRO J1655–40: “Ten epochs of 22.5-GHz images of the resolved radio source, from Aug. 18 to Sept. 5, show a linear expansion of two components at an angular rate of 0′′.056 ± 0′′.012 (3σ) per day on a line with a position angle of 54^{o}, with expansion beginning Aug. 12.6 UT. If the distance is 3.5 kpc (IAUC 6062), this expansion rate corresponds to a separation velocity of 1.1c perpendicular to the line of sight. Preliminary indications are that the southwestern component is stable at a position α = 16^{h}54^{m}00^{s}.14 ± 0^{s}.02, δ = −39^{o}50′44′′.9 ± 0′′.9 (equinox 2000.0), and the apparent motion is mostly from the northeastern component moving away from this position. Because of this behavior, astrometric work on the optical and radio positions of this object is urgently needed. On Sept. 5.2, the radio source’s intensity was 0.3, 0.15, 0.13, 0.11, and 0.09 Jy at 1.49, 4.9, 8.4, 14.9, and 22.5 GHz, respectively. In addition to the general trend of a flux decay with an e-folding time of ∼ 6 days, at frequencies below 5 GHz, there was an additional ∼ 0.3 Jy component in the light curves at higher frequencies that appeared ∼ Aug. 27, and which was present only in the southwestern component. There are also indications of variability on time scales of tens of minutes to hours in the same component. Part of the southwest component seems to be behaving like the fluctuating, slowly-decaying radio sources seen after the transient events in GRS 2023+338 and GRO J0422+32, and may be synchrotron radio emission in the outer portions of accretion-disk outflows.”

AG DRACONIS K. Petrik and L. Hric provide the following photometry obtained at the

Hlohovec Public Observatory, Slovakia (comparison stars BD +67^{o}925, HD 145991; estimated errors to last digit given parenthetically): July 12.05 UT, V = 8.55(3), B−V = +0.45(5), U −B = −1.08(6); 14.97, 8.52(2), +0.41(4), −0.81(9); 15.95, 8.55(1), +0.24(3), −0.94(6); 16.01, 8.51(1), +0.44(4), −1.23(4); 16.95, 8.50(2), +0.59(3), −1.25(1); 16.97, 8.57(2), +0.27(3), −1.23(4); Aug. 9.89, 8.68(2), +0.47(3), −0.85(6); 9.93, 8.63(1), +0.51(3), −0.96(4); 15.87, 8.75(1), +0.49(2), −1.15(3).

1994 September 8 Daniel W. E. Green